# Proca stars in excited states

### v1

@@ Abstract
In this paper we consider families of solutions for excited states of Proca stars in spherical symmetry. We focus on the first two excited configurations and perform a series of fully non-linear dynamical simulations in order to study their properties and stability. Our analysis reveals that excited Proca stars are always unstable against even very small perturbations, and their dynamical evolution can lead to three different final states: collapse to a black hole, dissipation, or migration to a different configuration in the ground state. We find that migration to the ground state can only occur in a small region of the parameter space of solutions with negative binding energy.
@@ I Introduction
Exotic stars are hypothetical self-gravitating compact objects composed of different types of exotic matter. One particular type of exotic compact object (ECO) are the bosonic solitons, which represent stationary solutions of the Einstein field equations minimally coupled to massive, scalar or vector, complex fields. The scalar case corresponds to the usual boson stars [1, 2], while the vector case corresponds to the so-called Proca stars [3]; these models are commonly referred to as bosonic stars. In addition to bosonic stars, there are also Dirac stars [4] composed of complex spinor fields, further expanding the array of potential ECOs proposed in the literature. For a comparative analysis and review of these types of objects see for example [5, 6].
ECOs are theorized to be detectable due to their gravitational interactions with surrounding matter. This detection has the potential to offer valuable insights into the nature of the dark sector of the Universe, in particular serving as a promising environment for identifying the nature of dark matter (DM). Indeed, bosonic fields are considered viable candidates for dark matter due to their ability to clump together and form compact objects. Depending on the value of the mass parameter of the bosonic field, these objects can be considered as massive compact objects in galactic halos (MACHO’s), or can even acquire galactic scales in the case of ultralight masses [7, 8, 9]. These objects can also exhibit characteristics that mimic black holes [10], serve as potential sources of dark matter, and manifest in a variety of other phenomena [11]. One extensively studied scenario involves investigating whether these bosonic fields could undergo accretion inside stars, leading to the formation of stable configurations with observable effects. For a more detailed discussion we direct the reader to [12].
@@
In this work we shall focus on Proca stars, that is spherically-symmetric self-gravitating configurations of a complex massive vector field (the Proca field) in four spacetime dimensions, and carrying a conserved Noether charge Q associated with a global U(1) symmetry. The first such self-gravitating solutions composed of massive spin-1 particles were reported in [3]. Numerical dynamical evolutions of these configurations have been performed in [13, 14], studying their stability and the various different phenomena that occur under perturbations of the Proca field in the ground state. Additionally, Proca stars have been set up in different astrophysical scenarios, such as those with rotation [15], charge [16], or more generalized models [17]. The primary objective of this work is to shed new light on the properties Proca stars in excited states, and in particular their stability under perturbations.
@@
This paper is organized as follows. In Section II we present the Einstein–Proca system. In Section III we derive the spherically symmetric field equations for the stationary configurations corresponding to Proca stars and also discuss the total mass, total bosonic charge, and binding energy of the system, as well as the boundary conditions and the numerical methods that we have used. Section IV presents our results for the families of stationary configurations in the first two excited states, as well as the resuts of their dynamical evolutions under perturbations. We conclude in Section V.
@@ II.1 Action and fields equations
with $G_{\mu\nu}$ the Einstein tensor for our curved spacetime and $T_{\mu\nu}$ is the stress-energy tensor associated with the Proca field, which be derived from the Lagrangian (1) and takes the form:
@@ II.2 The Proca equations in 3+1 formalism
where $\gamma_{ij}$ is the three dimensional metric defined within the spatial hypersurfaces, $\alpha$ is the lapse function, and $\beta^{i}$ the shift vector. The unit normal vector $n^{\mu}$ to the spatial hypersurfaces has components given by
@@
$W^{*}_{\mu\nu}:=-\frac{1}{2}\>e_{\alpha\beta\mu\nu}W^{\mu\nu}\;,$ (14)
with $e^{\alpha\beta\mu\nu}$ the Levi–Civita tensor. In terms of the electric and magnetic fields, the Proca field tensor and its dual can now be expressed as:
@@
$W^{*}_{\mu\nu}$ $=n_{\mu}{\cal B}_{\nu}-n_{\nu}{\cal B}_{\mu}-E_{\mu\nu\lambda}{\cal E}^{\lambda}\;.$ (16)
@@
with $d/dt:=\partial_{t}-\pounds_{\vec{\beta}}$, and where $D_{i}$ is the three-dimensional covariant derivative compatible with the spatial metric $\gamma_{ij}$ and $K$ is the trace of the extrinsic curvature tensor $K_{ij}$. Notice that again we only consider the spatial components of the electric and magnetic fields.
@@
One can also express the 3+1 matter terms coming from the stress–energy tensor (4) in terms of the 3+1 quantities associated with the Proca field. For the energy density we find:
@@
where now $e^{ijk}$ is the three dimensional Levi–Civita tensor defined as $e^{ijk}:=n_{\alpha}e^{\alpha ijk}$, and where c.c. indicates the complex conjugate of the previous terms.
Finally, for the 3D stress tensor we have:
@@ III Proca stars
with $\psi$ a conformal factor, and where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ is the standard solid angle element. In general the metric functions $(\alpha,\psi,A,B)$ are functions of both time $t$ and radius $r$, though for the case of stationary configurations they will depend only on $r$. Notice that in principle we don’t need to introduce the conformal factor $\psi$, and indeed we will take it to be equal to 1 for our stationary configurations, but we do need it for the BSSN formalism used in the dynamical simulations that we will consider later. In the context of this work, the assumption that the shift vector vanishes, $\beta^{i}=0$, is adequate for our time evolutions. In order to obtain our stationary configurations we will further restrict ourselves to using the areal radial coordinate, so in this Section we will take $B=\psi=1$.
Proca stars (PSs) are self-gravitating solutions in spherical symmetry for which the Proca field exhibits a harmonic time dependence of the form of $e^{-i\omega t}$, resulting in a time-independent stress-energy tensor which leads to a static spacetime metric. Notice that in spherical symmetry the magnetic field vanishes, while the vector potential and the electric field only have a non-zero radial component. We then propose the following ansatz for the scalar potential and the radial components of the vector potential and the electric field:
@@
with $\omega$ a real constant corresponding to the oscillation frequency of the Proca field, and where the functions $\varphi(r)$, $a(r)$ and $e(r)$ are real and depend only on the radial coordinate $r$. The explicit factor of $i$ introduced in the ansatz for the radial component of the vector potential $a_{r}$ is there in order to guarantee that both the momentum density (25) and the particle flux coming from the conserved Noether current (28) vanish, as they should for a static solution. In our case the momentum density reduces to:
@@
By substituting our harmonic anzats (30) into the evolution equations (17)—(19) and the Gauss constraint (21), we can derive a set of ordinary differential equations in $r$ to solve for our Proca star configurations. In order to simplify some expressions we work with $F(r):=\alpha\varphi$ instead of $\varphi$ itself. Once the system is solved, $\varphi$ can be immediatly reconstructed as $\varphi=F/\alpha$.
@@
$\partial_{r}A=A\left[\frac{(1-A)}{r}+8\pi rA\rho\right]\;,$ (38)
with $\rho$ the energy density. The equation for the lapse comes from our choice of using the areal radius, which in turn implies that the angular component of the extrinsic curvature has to vanish for all time, which through the ADM equations leads to the so-called “polar-areal” gauge condition that takes the form:
@@
Notice that the equations for $\alpha$ and $A$ above have matter terms on the right hand side corresponding to the energy density $\rho$ and the component $S^{r}_{r}$ of the stress tensor. Using now equations (23) and (26) one can show that in our case those quantities reduce to:
@@
As a final comment, equation (36) above has terms on the right hand side that involve derivatives of $\alpha$ and $A$. But we can substitute those derivatives using the equations for $\alpha$ and $A$ that we just found, so that the equation for $a$ now becomes:
@@ III.1 Boundary conditions
We need to solve the system of first order differential equations (38), (39), (35), (42) for the functions $(A,\alpha,F,a)$ in order to obtain the family of Proca stars, subject to a series of physical boundary conditions. The boundary conditions for the radial metric come from demanding regularity at the origin:
@@
From this we see that in order for $\varphi$ to be real we must necessarily have $\omega^{2}<\alpha^{2}_{\infty}m^{2}$. But notice that in that case we can have both increasing and decreasing exponential solutions. The increasing solutions are unphysical, as they are not compatible with an asymptotically flat spacetime. In fact, one finds that we only have exponentially decaying solutions for a discrete set of values of the frequency $\omega$, which implies that we need to solve an eigenvalue problem.
@@
with $\alpha_{0}=\alpha(r=0)$ and $\varphi_{0}=\varphi(r=0)$. In the same way we see that in the equation for $\partial_{r}A$ there is a term that goes as $(1-A)/r$. This means that close to the origin we must have $A\sim 1+k_{A}r^{2}$. Substituting this in the equation we now find:
@@
Numerically, we solve our system of equations starting from the origin, where we choose as our free parameter the value of $\varphi_{0}$ and a trial value for the frequency $\omega$. We then integrate the system outwards (using a fourth order Runge–Kutta method), and keep modifying the value of $\omega$ with a shooting algorithm until we find a solution that decays far away.
There is one last detail that must be considered. Since initially we don’t know the value of the lapse $\alpha$ at the origin we start by taking $\alpha_{0}=1$. However, what we really want is $\alpha$ to become zero at infinity. But this is not a serious problem since the equation for $\partial_{r}\alpha$ is linear in $\alpha$, so we can always rescale the solution at the end using the asymptotic value of the lapse $\alpha_{\infty}$ (obtained by extrapolation assuming that the lapse far away behaves as $\alpha\simeq\alpha_{\infty}-c/r$ for some positive constant $c$). However, in order for this to still be a solution of the whole system we must also rescale the frequency $\omega$ by the same factor. In particular this implies that our initial trial frequency is in general such that $\omega>m$, but once we rescale it we will have $\omega<m$. Notice also that if we solve the system for $F$ instead of $\varphi$, then we also need to rescale $F$ by the same factor in order to maintain $\varphi=F/\alpha$ the same as before.
@@ III.2 Total mass and particle number
in order to isolate the kinetic and potential contributions. This means that those solutions with $mQ>M$ will have a negative binding energy, $U<0$, and will correspond to gravitationally bound stars. Conversely, solutions with $mQ<M$ have positive binding energy $U>0$, making them gravitationally unbound (and hence should be expected to be unstable against perturbations).
We can also use the mass integral above to define an effective radius $R_{99}$ which contains 99% of the total mass. This effective radius is necessary due to the fact that, as we mentioned above, for Proca stars the energy density decays exponentially far away implying that there is no real surface.
@@ III.3 Perturbations
In previous sections we derived the equations that we need to solve in order to obtain stationary Proca star configurations, we will show families of such solutions for the ground state and the first two excited states in Section IV.1 below. However, one of the aims of this paper is to study the dynamical stability of perturbed Proca star configurations. We will therefore also consider self-consistent perturbations to the stationary Proca star solutions.
Once we have a stationary solution we will construct a perturbation to it in the following way:
@@
    We add a small perturbation to the scalar potential $\varphi(r)$ (in practice we add it to $F(r)$), typically a gaussian, leaving the vector potential $a(r)$ unchanged. The perturbation in $\varphi$ (or $F$) must be even with respect to reflections on the origin in order to maintain regularity there. For example, if the add a gaussian centered on a point $r=r_{0}$ away from the origin, we must also add a similar gaussian centered on $r=-r_{0}$ in order to guarantee that we still have $\partial_{r}\varphi(r=0)=0$. Notice also that if we only modify the real part of the scalar potential while leaving the vector potential purely imaginary, the momentum density $J_{r}$ remains zero. Similarly, since the scalar potential remains real the electric field $\mathcal{E}^{r}$ will also be purely real, so that the particle flux $j^{r}_{Q}$ also remains zero.
@@
The procedure just described results in consistent initial data for a perturbed Proca star. Of course, our perturbed star will no longer be stationary. However, as mentioned above, the momentum density and particle flux are initially zero, which is consistent with taking an initially vanishing extrinsic curvature. Our perturbed initial data will then result in a momentarily static geometry corresponding to a moment of time symmetry.
Once we have the perturbed initial data we evolve it in time using the OllinSphere code already described in [22, 23, 24, 25, 26], which implements a BSSN formulation adapted to spherical symmetry [21]. For the evolution of the Proca field variables we use equations (17)-(19), which in our case reduce to:
@@
Notice that even if for our initial data we took the conformal factor $\psi$ and the angular metric component $B$ equal to 1 in the metric (29), this will not remain true for a dynamical simulation. One must also remember that all three quantities $(\phi,a_{r},\mathcal{E}^{r})$ are complex, so we must evolve their real and imaginary parts separately.

### v2

@@ Abstract
In this paper, we consider families of solutions for excited states of Proca stars in spherical symmetry. We focus on the first two excited configurations and perform a series of fully non-linear dynamical simulations in order to study their properties and stability. Our analysis reveals that excited Proca stars are always unstable against even very small perturbations, and their dynamical evolution can lead to three different final states: collapse to a black hole, dissipation, or migration to a different configuration in the ground state. We find that migration to the ground state can only occur in a small region of the parameter space of solutions with negative binding energy.
@@ I Introduction
Exotic stars are hypothetical self-gravitating compact objects composed of different types of exotic matter. One particular type of exotic compact object (ECO) is the bosonic soliton, which represents stationary solutions of the Einstein field equations minimally coupled to massive scalar or vector complex fields. The scalar case corresponds to the usual boson stars [1, 2], while the vector case corresponds to the so-called Proca stars [3]; these models are commonly referred to as bosonic stars.
There exists a variety of subtypes of bosonic star models that include self-interaction terms, which provide additional parameters to better resemble astrophysical systems; see [4, 5] for reviews. In addition to bosonic stars, there are also oscillatons [6] composed of real scalar fields, and Dirac stars [7] composed of complex spinor fields, further expanding the array of potential ECOs proposed in the literature. For a comparative analysis and review of these types of objects, see [8, 9].
ECOs are theorized to be detectable because of their gravitational interactions with surrounding matter. Extensive research into these compact objects has explored their physical properties, stability, and dynamical behaviors, making them promising candidates for the dark matter component of the Universe. Their ability to clump together and form compact structures, combined with their versatility shaped by the choice of different matter fields and self-interaction potentials, allows them to mimic a variety of astrophysical objects and phenomena. Recently, bosonic stars have been modeled in various astrophysical contexts, such as bosonic dark matter in galactic halos to address the rotation curve problem for galaxies [10]. Other studies have analyzed lensing imagescreated by shadows cast in the presence of accretion disks [11, 12], as well as simulations of vector/boson star mergers and the prediction of their gravitational wave signals in order to compare them with the LIGO/Virgo catalog [13, 14]. For a more detailed discussion, we direct the reader to references [15, 16]. All this research offers valuable insights into the nature of the dark sector of the Universe.
@@
In this work, we shall focus on Proca stars, that is, self-gravitating configurations of a complex massive vector field (the Proca field) in four spacetime dimensions, and carrying a conserved Noether charge associated with a global U(1) symmetry. The first such self-gravitating solutions composed of massive spin-1 particles were reported in [3]. Numerical dynamical evolutions of these configurations have been performed in [17, 18], studying their stability and the various different phenomena that occur under perturbations of the Proca field in the ground state. Additionally, Proca stars have been set up in different astrophysical scenarios, such as those with rotation [19], charge [20], or more generalized models [21].
In this paper, we focus on excited states of Proca stars, which are significant because they may represent intermediate states during the formation of ECOs. These excited states exhibit distinct properties when compared to the ground state, and have been extensively studied in the case of scalar boson stars in both spherical and axial symmetry [22, 23, 24]. The primary objective of our work is to extend these studies to the case of excited Proca stars, concentrating particularly on their stability under perturbations.
We should emphasize that in this work we will only consider the spherically symmetric case. In particular, the ground state will be defined as that spherical configuration for which the radial component of the Proca vector potential has no nodes, or alternatively as the spherical state with minimum total mass for a given central value of the Proca scalar potential, while excited states will correspond to solutions with nodes (see sections below for details). It is important to mention here that it was recently shown in [25] that this spherically symmetric “ground state” is not stable against non-spherical perturbations, and it does correspond to the minimum mass configuration which is in fact non-spherical. However, here we will only consider spherically symmetric configurations and perturbations.
@@
This paper is organized as follows. In Section II we present the Einstein–Proca system. In Section III we derive the spherically symmetric field equations for the stationary configurations corresponding to Proca stars and also discuss the total mass, total bosonic charge, and binding energy of the system, as well as the boundary conditions and the numerical methods that we have used. Section IV presents our results for the families of stationary configurations in the first two excited states, as well as the results of their dynamical evolutions under perturbations. We conclude in Section V.
@@ II.1 Action and fields equations
where $G_{\mu\nu}$ is the Einstein tensor for our curved spacetime and $T_{\mu\nu}$ the stress-energy tensor associated with the Proca field, which is derived from the Lagrangian (1) and takes the form:
@@ II.2 The Proca equations in 3+1 formalism
where $\gamma_{ij}$ is the three dimensional metric defined within the spatial hypersurfaces, $\alpha$ is the lapse function, and $\beta^{i}$ is the shift vector. The unit normal vector $n^{\mu}$ to the spatial hypersurfaces has components given by
@@
$W^{*}_{\alpha\beta}:=-\frac{1}{2}\>E_{\alpha\beta\mu\nu}W^{\mu\nu}\;,$ (14)
with $E_{\alpha\beta\mu\nu}$ the Levi–Civita tensor defined in terms of the totally anti-symmetric symbol $\epsilon_{\alpha\beta\mu\nu}$ as $E_{\alpha\beta\mu\nu}:=\left|g\right|^{1/2}\epsilon_{\alpha\beta\mu\nu}$, with $g$ the determinant of the spacetime metric $g_{\mu\nu}$ (we use a convention such that $\epsilon_{0123}=1=-\epsilon^{0123}$). Notice that from the above definitions one can show that the electric and magnetic fields are purely spatial, that is, $n^{\mu}\mathcal{E}_{\mu}=n^{\mu}\mathcal{B}_{\mu}=0$.
In terms of the electric and magnetic fields, the Proca field tensor and its dual can now be expressed as:
@@
$W^{*}_{\mu\nu}$ $=n_{\mu}{\cal B}_{\nu}-n_{\nu}{\cal B}_{\mu}-E_{\mu\nu\lambda}{\cal E}^{\lambda}\;,$ (16)
where the three-dimensional Levi–Civita tensor is now defined as $E_{\mu\nu\lambda}:=n^{\alpha}E_{\alpha\mu\nu\lambda}$.
@@
with $d/dt:=\partial_{t}-\pounds_{\vec{\beta}}$ with £$\pounds_{\vec{\beta}}$ the Lie derivative with respect to the shift vector, and where $D_{i}$ is the three-dimensional covariant derivative compatible with the spatial metric $\gamma_{ij}$, and $K$ is the trace of the extrinsic curvature tensor $K_{ij}$. Notice that again we only consider the spatial components of the electric and magnetic fields.
@@
One can express the 3+1 matter terms coming from the stress–energy tensor (4) in terms of the 3+1 quantities associated with the Proca field. For the energy density we find:
@@
where c.c. indicates the complex conjugate of the previous terms. Finally, for the 3D stress tensor we have:
@@ III Proca stars
with $\psi$ a conformal factor, and where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ is the standard solid angle element. In general, the metric functions $(\alpha,\psi,A,B)$ are functions of both time $t$ and radius $r$, though for the case of stationary configurations they will depend only on $r$. Notice that in principle we do not need to introduce the conformal factor $\psi$, and indeed we will take it to be equal to 1 for our stationary configurations, but we do need it for the BSSN formalism used in the dynamical simulations that we will consider later. In the context of this work, the assumption that the shift vector vanishes, $\beta^{i}=0$, is adequate for both stationary configurations and our dynamical evolutions. In order to obtain our stationary configurations, we will further restrict ourselves to using the areal radial coordinate, so in this section we will take $B=\psi=1$.
Proca stars (PSs) are self-gravitating solutions for which the Proca field exhibits a harmonic time dependence, resulting in a time-independent stress-energy tensor which leads to a static spacetime metric. As mentioned above, here we will only consider the case of spherical symmetry. Notice that in spherical symmetry the magnetic field vanishes, while the vector potential and the electric field only have a non-zero radial component. We then propose the following ansatz for the scalar potential and the radial components of the vector potential and the electric field:
@@
with $\omega$ a real constant corresponding to the oscillation frequency of the Proca field, and where the functions $\varphi(r)$, $a(r)$ and $e(r)$ are real and depend only on the radial coordinate $r$. The explicit factor of $i$ introduced in the ansatz for the radial component of the vector potential $a_{r}$ is there to guarantee that both the momentum density (25) and the particle flux coming from the conserved Noether current (28) vanish, as they should for a static solution. In our case the momentum density reduces to:
@@
By substituting our harmonic anzats (30) into the evolution equations (17)—(19) and the Gauss constraint (21), we can derive a set of ordinary differential equations in $r$ to solve for our Proca star configurations. In order to simplify some expressions, we work with $F(r):=\alpha\varphi$ instead of $\varphi$ itself. Once the system is solved, $\varphi$ can be immediately reconstructed as $\varphi=F/\alpha$.
@@
$\partial_{r}A=A\left[\frac{(1-A)}{r}+8\pi rA\rho\right]\;.$ (38)
The equation for the lapse comes from our choice of using the areal radius, which in turn implies that the angular component of the extrinsic curvature has to vanish for all time, which through the ADM equations leads to the so-called “polar-areal” gauge condition that takes the form:
@@
Notice that the equations for $\alpha$ and $A$ above have matter terms on the right-hand side corresponding to the energy density $\rho$ and the component $S^{r}_{r}$ of the stress tensor. Using now equations (23) and (26) one can show that in our case those quantities reduce to:
@@
As a final comment, equation (36) above has terms on the right-hand side that involve derivatives of $\alpha$ and $A$. But we can substitute those derivatives using the equations for $\alpha$ and $A$ that we just found, so that the equation for $a$ now becomes:
@@ III.1 Boundary conditions
We need to solve the system of first-order differential equations (35), (38), (39), (42) for the functions $(A,\alpha,F,a)$ in order to obtain the family of Proca stars, subject to a series of physical boundary conditions. The boundary conditions for the radial metric come from demanding regularity at the origin:
@@
From this we see that in order for $\varphi$ to be real we must necessarily have $\omega^{2}<\alpha^{2}_{\infty}m^{2}$. But notice that in that case we can have both increasing and decreasing exponential solutions. The increasing solutions are unphysical, as they are not compatible with an asymptotically flat spacetime. In fact, one finds that we only have exponentially decaying solutions for a discrete set of values for the frequency $\omega$, which implies that we need to solve an eigenvalue problem.
@@
with $\alpha_{0}=\alpha(r=0)$ and $\varphi_{0}=\varphi(r=0)$. In the same way, we see that in the equation for $\partial_{r}A$ there is a term that goes as $(1-A)/r$. This means that close to the origin we must have $A\sim 1+k_{A}r^{2}$. Substituting this in the equation we now find:
@@
Numerically, we solve our system of equations starting from the origin, where we choose as our free parameter the value of $\varphi_{0}$ and a trial value for the frequency $\omega$. We then integrate the system outward (using a fourth-order Runge–Kutta method) and keep modifying the value of $\omega$ with a shooting algorithm until we find a solution that decays far away.
There is one last detail that must be considered. Since initially we do not know the value of the lapse $\alpha$ at the origin we start by taking $\alpha_{0}=1$. However, what we really want is $\alpha$ to become zero at infinity. But this is not a serious problem since the equation for $\partial_{r}\alpha$ is linear in $\alpha$, so we can always rescale the solution at the end using the asymptotic value of the lapse $\alpha_{\infty}$ (obtained by extrapolation assuming that the lapse far away behaves as $\alpha\simeq\alpha_{\infty}-c/r$ for some positive constant $c$). However, in order for this to still be a solution of the whole system we must also rescale the frequency $\omega$ by the same factor. In particular, this implies that our initial trial frequency is in general such that $\omega>m$, but once we rescale it we will have $\omega<m$. Notice also that if we solve the system for $F$ instead of $\varphi$, then we also need to rescale $F$ by the same factor to maintain $\varphi=F/\alpha$ the same as before.
@@ III.2 Total mass and particle number
in order to isolate the kinetic and potential contributions. This means that those solutions with $mQ>M$ will have a negative binding energy $U<0$ and will correspond to gravitationally bound stars. Conversely, solutions with $mQ<M$ have positive binding energy $U>0$, making them gravitationally unbound. In the case of scalar boson stars gravitationally unbound configurations have been found to be unstable against perturbations [30, 31, 32], so one might expect the same to be true for Proca stars, and this is indeed what we find in our simulations (see below). The definition of the binding energy above has been used frequently in the literature on bosonic stars, starting from the early work on boson stars (see e.g. [33]). A formal justification for this definition can be found in [34].
We can also use the mass integral above to define an effective radius $R_{99}$, which contains 99% of the total mass. This effective radius is necessary because, as mentioned above, for Proca stars the energy density decays exponentially implying that there is no real surface.
@@ III.3 Perturbations
In previous sections we derived the equations that we need to solve in order to obtain stationary Proca star configurations; we will show families of such solutions for the ground state and the first two excited states in Section IV.1 below. However, one of the aims of this paper is to study the dynamical stability of perturbed Proca star configurations. We will therefore also consider self-consistent perturbations to the stationary Proca star solutions.
Once we have a stationary solution, we will construct a perturbation to it in the following way:
@@
    We add a small perturbation to the scalar potential $\varphi(r)$ (in practice, we add it to $F(r)$), typically a Gaussian, leaving the vector potential $a(r)$ unchanged. The perturbation in $\varphi$ (or $F$) must be even with respect to reflections on the origin to maintain regularity there. For example, if we add a Gaussian centered on a point $r=r_{0}$ away from the origin, we must also add a similar Gaussian centered on $r=-r_{0}$ to guarantee that we still have $\partial_{r}\varphi(r=0)=0$. Notice also that if we only modify the real part of the scalar potential while leaving the vector potential purely imaginary, the momentum density $J_{r}$ remains zero. Similarly, since the scalar potential remains real, the electric field $\mathcal{E}^{r}$ will also be purely real, so that the particle flux $j^{r}_{Q}$ also remains zero.
@@
The procedure just described results in consistent initial data for a perturbed Proca star. Of course, the perturbed star will no longer be stationary. However, as mentioned above, the momentum density and particle flux are initially zero, which is consistent with taking an initially vanishing extrinsic curvature. Our perturbed initial data will then correspond to a moment of time symmetry, that is, the initial time derivative of the spatial metric will vanish, but its second time derivative will not.
Once we have the perturbed initial data we evolve it in time using the OllinSphere code (see below), which implements a BSSN formulation adapted to spherical symmetry [29]. For the evolution of the Proca field variables we use equations (17)-(19), which in our case reduce to:
@@
Notice that even if for our initial data we took the conformal factor $\psi$ and the angular metric component $B$ equal to 1 in the metric (29), this will not remain true for a dynamical simulation. We must also remember that all three quantities $(\phi,a_{r},\mathcal{E}^{r})$ are complex, so we must evolve their real and imaginary parts separately.
Finally, as gauge conditions during the evolution we take a vanishing shift vector $\beta^{r}=0$, and we use the well-known “1+log” slicing condition which takes the form:
$\partial_{t}\alpha=-2\alpha K\;.$ (59)
This slicing condition has been shown to be very robust in practice, allowing for stable and long lived dynamical evolutions [26, 35, 36, 37].